OpenStudy (anonymous):

The length of satisfactory service (years) provided by a certain model of laptop computers is a random variable having the probability density: f(x)=1/4e^-x/4.5 for x>0 0 for x<= 0 Find the probabilities that one of these laptops will provide satisfactory service for (a) at most 2.5 years. (b)anywhere from 4 to 6 years (c) at least 6.75

OpenStudy (anonymous):

So for part (a) from guessing that I have to convert the function from f(x) to a distribution function F(x) and then plug in 2.5 for F(x) to get my answer.

OpenStudy (anonymous):

So far I have: \[F(x)=\int\limits_{0}^{x}\frac{ 1 }{ 4.5 } e ^{-x/4.5}=.2222e^{-.2222x}dx\]

OpenStudy (anonymous):

So F(2.5)= \[.2222e^{-.2222*(2.5)}=-.5738\]

OpenStudy (anonymous):

nevermind question solved.